# Decomposing Disease Descriptions for Enhanced Pathology Detection: A Multi-Aspect Vision-Language Matching Framework

### v1

@@ Abstract
Medical vision language pre-training (VLP) has emerged as a frontier of research, enabling zero-shot pathological recognition by comparing the query image with the textual descriptions for each disease. Due to the complex semantics of biomedical texts, current methods struggle to align medical images with key pathological findings in unstructured reports. This leads to the misalignment with the target disease’s textual representation. In this paper, we introduce a novel VLP framework designed to dissect disease descriptions into their fundamental aspects, leveraging prior knowledge about the visual manifestations of pathologies. This is achieved by consulting a large language model and medical experts. Integrating a Transformer module, our approach aligns an input image with the diverse elements of a disease, generating aspect-centric image representations. By consolidating the matches from each aspect, we improve the compatibility between an image and its associated disease. Additionally, capitalizing on the aspect-oriented representations, we present a dual-head Transformer tailored to process known and unknown diseases, optimizing the comprehensive detection efficacy. Conducting experiments on seven downstream datasets, ours outperforms recent methods by up to 8.07% and 11.23% in AUC scores for seen and novel categories, respectively. Our code is released at https://github.com/HieuPhan33/MAVL.
@@ 3.1 Problem setting
3.2 Multi-aspect vision-language matching for enhanced pathology detection
@@
Figure 4: Multi-aspect vision-language matching framework (MAVL) decomposes diseases into a set of shared visual aspects using LLM and prior knowledge from two medical experts. An aspect-oriented dual-head Transformer (A) visually searches for the queried visual aspects in the image and maximizes detection accuracy of both unseen and seen diseases via two learning signals. The contrastive head (B) captures generalizable features and performs zero-shot classification of unseen diseases by comparing visual aspects between the target disease and the healthy category. The supervised head (C) captures discriminative features to classify fine-grained seen diseases.
@@
By reformulating VLP as a multi-label classification (Sec. 3.1), we directly train the visual aspect-query Transformer using supervision labels $Y$, thus improving zero-shot classification of seen diseases on downstream tasks. The supervised head is trained via a cross entropy loss $\mathcal{L}_{\text{sup}}=\text{CE}(\hat{p},Y)$. Besides, we predict the location embeddings $\tilde{e}_{j}$ of each disease from the its features $\mathcal{F}_{j}$ using a projection layer $W_{\text{loc}}$. We adopt the location contrastive loss $\mathcal{L}_{\text{loc}}$ from [39] to match the predicted location embedding $\tilde{e}_{j}$ of a positive disease with the embedding of ground-truth location $e_{j}=f_{\text{text}}(\text{``It is located at }\{l_{j}\}\text{"})$ using embedding $\bar{e}$ of other $L_{\text{neg}}=L\setminus\{l_{j}\}$ as negative samples. The network is trained end-to-end with the combined loss objectives:
@@ 5.3 Ablation study
Effectiveness of multi-aspect decomposition for VLP. This section answers two questions: (i) whether adding more concepts improves zero-shot performance, and (ii) which concept matching yields higher performance gain. Fig. 6(a) presents the performance when adding a single aspect. Matching with any single aspect boosts the accuracy of matching with a disease’s definition as in [39]. Notably, the gains on on unseen PadChest and Covid CXR-2 are 3.64% and 4.02%, respectively when using the aspect ‘texture’. Distinct aspects such as texture, patterns, and shapes yield higher gains. Fig. 6(b) shows the AUC scores when adding increasing numbers of visual aspects. Matching with higher numbers of aspects improves the zero-shot disease classification.
@@ 6 Conclusion
This study introduces a Multi-Aspect Vision-Language Matching (MAVL) framework, designed to intricately dissect pathological entities into distinct, visually descriptive aspects, thereby significantly enhancing pathology detection. The MAVL framework incorporates a dual-head, aspect-oriented Transformer, which decouples learning signals to optimize performance across both seen and unseen diseases. This approach is particularly effective for novel diseases, achieving a remarkable 11.23% performance gain over existing methods. Moreover, MAVL’s aspect-based matching leads to an impressive 18.46% AUC improvement for rare unseen diseases.
Acknowledgement. This research was partially funded by the Hospital Research Foundation Group and supported by Dr. Perperidis from Adelaide’s Women’s and Children’s Hospital Network to raise necessary funding.

### v2

# Decomposing Disease Descriptions for Enhanced Pathology Detection: A Multi-Aspect Vision-Language Pre-training Framework

@@ Abstract
Medical vision language pre-training (VLP) has emerged as a frontier of research, enabling zero-shot pathological recognition by comparing the query image with the textual descriptions for each disease. Due to the complex semantics of biomedical texts, current methods struggle to align medical images with key pathological findings in unstructured reports. This leads to the misalignment with the target disease’s textual representation. In this paper, we introduce a novel VLP framework designed to dissect disease descriptions into their fundamental aspects, leveraging prior knowledge about the visual manifestations of pathologies. This is achieved by consulting a large language model and medical experts. Integrating a Transformer module, our approach aligns an input image with the diverse elements of a disease, generating aspect-centric image representations. By consolidating the matches from each aspect, we improve the compatibility between an image and its associated disease. Additionally, capitalizing on the aspect-oriented representations, we present a dual-head Transformer tailored to process known and unknown diseases, optimizing the comprehensive detection efficacy. Conducting experiments on seven downstream datasets, ours improves the accuracy of recent methods by up to 8.56% and 17.0% for seen and unseen categories, respectively. Our code is released at https://github.com/HieuPhan33/MAVL.
@@ 3.1 Problem setting
3.2 Multi-aspect vision-language pre-training for enhanced pathology detection
@@
Figure 4: Multi-aspect vision-language pre-training framework (MAVL) decomposes diseases into a set of shared visual aspects using LLM and prior knowledge from two medical experts. An aspect-oriented dual-head Transformer (A) visually searches for the queried visual aspects in the image and maximizes detection accuracy of both unseen and seen diseases via two learning signals. The contrastive head (B) captures generalizable features and performs zero-shot classification of unseen diseases by comparing visual aspects between the target disease and the healthy category. The supervised head (C) captures discriminative features to classify fine-grained seen diseases.
@@
By reformulating VLP as a multi-label classification (Sec. 3.1), we directly train the visual aspect-query Transformer using supervision labels $Y$, thus improving zero-shot classification of seen diseases on downstream tasks. The supervised head is trained via a cross entropy loss $\mathcal{L}_{\text{sup}}=\text{CE}(\hat{p},Y)$. Besides, the network predicts the location embedding $\tilde{e}_{j}$ of each disease from its features $\mathcal{F}_{j}$ using a simple projection layer. If the disease $c_{j}$ is positively present and appears at location $l_{j}$, we adopt the location contrastive loss $\mathcal{L}_{\text{loc}}$ from [39] to match the predicted location embedding $\tilde{e}_{j}$ with the positive location embedding $e_{j}=f_{\text{text}}(\text{``It is located at }\{l_{j}\}\text{"})$. The network is trained end-to-end with the combined loss objectives:
@@ 5.3 Ablation study
Effectiveness of multi-aspect decomposition for VLP. This section answers two questions: (i) whether adding more concepts improves zero-shot performance, and (ii) which concept matching yields higher performance gain. Fig. 6(a) presents the performance when adding a single aspect. Matching with any single aspect boosts the accuracy of matching with a disease’s definition as in [39]. Notably, the gains on on unseen PadChest and Covid CXR-2 are 3.64% and 4.02%, respectively when using the aspect ‘texture’. More distinct aspects such as texture, patterns, and shapes yield higher gains. Fig. 6(b) shows the AUC scores when adding increasing numbers of visual aspects. Matching with higher numbers of aspects improves the zero-shot classification.
@@ 6 Conclusion
This study introduces a Multi-Aspect Vision-Language Matching (MAVL) framework, designed to intricately dissect pathological entities into distinct, visually descriptive aspects, thereby significantly enhancing pathology detection. The MAVL framework incorporates a dual-head, aspect-oriented Transformer, which decouples learning signals to optimize performance across both seen and unseen diseases. This approach is particularly effective for novel diseases, achieving a remarkable 16.76% performance gain over existing methods. Moreover, MAVL’s aspect-based matching leads to an impressive 17.0% AUC improvement for rare unseen diseases.
Acknowledgement. This research was partially funded by the Hospital Research Foundation Group and supported by Dr. Perperidis from Adelaide’s Women’s and Children’s Health Network to raise necessary funding.